A group of researchers hypothesize that broken bones will take longer to heal among smokers versus nonsmokers. They test this question with a sample of 20 smokers, where the average time to heal was 32 days (SS = 1161), and a sample of 45 nonsmokers, where the average time to heal was 21 days (SS = 1179). Use the critical-value approach for the appropriate statistical test (where α = .05) to determine if there is a significant difference in bone healing time between smokers and nonsmokers. 1.List your estimated standard error 4. List your t test statistic (where: smokers – nonsmokers) 7. Calculate Cohen’s d for the difference in healing time
Answer #1

Ans:

1)pooled standard deviation=sqrt((1161+1179)/(20+45-2))=6.094

standard error =6.094*sqrt((1/20)+(1/45))=1.638

4)

Test statistic:

t=(32-21)/1.638

t=6.716

df=20+45-2=63

7)

Cohen 'd =(32-21)/6.094=1.81